From mar@roe.ac.uk Fri Oct 31 08:39:26 2014
Date: Fri, 31 Oct 2014 08:30:45 +0000
From: Mike Read <mar@roe.ac.uk>
To: VDFS Science Archive Team -- Clive Davenhall <acd@roe.ac.uk>,
     Dave Morris <dmr@roe.ac.uk>, Eckhard Sutorius <etws@roe.ac.uk>,
     Mark Holliman <msh@roe.ac.uk>, Nicholas Cross <njc@roe.ac.uk>,
     Nigel Hambly <nch@roe.ac.uk>, Rob Blake <rpb@roe.ac.uk>,
     Ross Collins <rsc@roe.ac.uk>, Stelios Voutsinas <stv@roe.ac.uk>
Cc: CCs for VDFS -- Alastair Edge <Alastair.Edge@durham.ac.uk>,
     Andrew Lawrence <al@roe.ac.uk>, Bob Mann <rgm@roe.ac.uk>,
     Dante Minniti <dante@astro.puc.cl>, James Dunlop <jsd@roe.ac.uk>,
     Jim Emerson <j.p.emerson@qmul.ac.uk>, Keith Noddle <keith@keithnoddle.org>,
     Lorenzo Rimoldini <Lorenzo.Rimoldini@unige.ch>,
     Maria-Rosa Cioni <mcioni@aip.de>, matt.jarvis@astro.ox.ac.uk,
     Mike Irwin <mike@ast.cam.ac.uk>,
     Nigel Metcalfe <nigel.metcalfe@durham.ac.uk>,
     Norman Gray <norman@astro.gla.ac.uk>, Phil Lucas <pwl@star.herts.ac.uk>,
     Richard McMahon at IOA <rgm@ast.cam.ac.uk>,
     Simon Dye <Simon.Dye@nottingham.ac.uk>,
     Stephen Warren <s.j.warren@ic.ac.uk>, Tom Kerr <tkerr@jach.hawaii.edu>,
     Tom Shanks <tom.shanks@durham.ac.uk>
Subject: WFAU VDFS Science Archives progress meeting minutes, 24 October 2014

WFAU VDFS Science Archives progress meeting minutes, 24 October 2014

Present: MAR, RPB, RSC, RGM, ACD, NJC
Apologies: STV, DM, NCH, KTN, AL, ETWS, MSH

DoNM: 10am Friday 7 November 2014 in the Villas meeting room,
   ACD in the chair (room is booked).


Actions discharged this time:
----------------------------

ACTION: ACD to review VVV brainstorming Twiki page and decide who to
chase about progressing their assigned actions.
Discharged, in real time during the meeting! The status of each task was 
discussed.
NEW ACTION: All to finish off VVV brainstorming tasks by end of January at the
latest.
NEW ACTION: RSC to contact Iraklis Klampanos to discuss performance of 
alternative RDBMSs?


ACTION: RPB/ETWS to see which data on /sahure were no longer required.
Discharged, ETWS, RPB and MAR had removed unwanted files.
NEW ACTION: MSH to re-configure RAID on /sahure.

ACTION: RPB to check that status of the VVV DR2 ESO release
   scheduled for 15/9/14.
Discharged. Phil Lucas was performing some final checks.


Actions carried forward from 3/10/2014 meeting:
----------------------------------------------

ACTION: RPB (was MAR) to have a look at the status of non-survey
     releases for registered projects.
MAR had compiled a list of non-survey projects that could be released
and sent it to RPB.  RPB had yet to make these data available.

ACTION: RGM to investigate the size and content (how does it differ
   from DR9?) of DR10 to see if it's worth us hosting and whether it
   is necessary to keep all of the other releases.
Continues.  The first part had been done: the difference between DR9
and DR10 is largely the presence of additional spectra.  However,
the action continues as RGM had not yet established which early
releases could safely be scrapped.  For further details see the
WFAU SDSS strategy wiki page:
  http://apache.roe.ac.uk/twiki/bin/view/WFAU/SDSSDRPlan.

ACTION: ETWS, with RPB, to develop an automated helper script to
     keep release DB files links updated.
Continues: this action still has low priority.

ACTION: ACD speak to HME about the possibility of IT support
   making nightly backups of the SVN repository.
Continues (HME on leave).

ACTION: RPB write a chron job to make daily check-outs of all the
   WFAU Projects in the SVN repository in order to provide a backup
   snapshot of the current state.
Continues.


Specific points and new actions:
-------------------------------

Project management:

RGM reported that extra money for hardware had been secured and that the grant
would be extended another year, up to April 2016. The bad news being that PDRs
and another WFAU quarterly meeting would take place before Christmas.


WFCAM, VISTA and VST updates:

RGM noted that the UKIRT handover would take place next week.


Review of Deadlines:

The list of recently-passed and forthcoming deadlines was reviewed
and the list will be updated to reflect this discussion.  The list
can be viewed at:

   http://apache.roe.ac.uk/twiki/bin/view/WFAU/ForthcomingDeadlines


Networking:

Assumed working (ETWS on leave).


WSA/VSA/OSA/GES Operations:

ACD noted that he'd copied over the GES iDR3 parameter files.

OSA CU4 ingests were running


Hardware and Systems:

RPB reported that the new tape library was working, though a new tape drive
as needed.

They'd been a few disk failures on sheshonq and pedubaste but these had been
fixed.


Software:

NJC noted that he'd been working on tweaks to the ESO release code, the 
matched aperture pipeline and proper motions.

RSC reported software updates to the curation code following the discovery
of problems arising from a STARLINK update to the AST library. Since the 
AST update a few months ago parts of the codebase had been incorrectly
doing pixel <=> sky coordinates conversions for TAN projections. This had 
affected a couple of VSA releases, these would be replaced or 
superseded shortly. A note had been placed on the known issues page.


Survey Data Release:

A GES release had been made, gesiDR4LineList. This release currently only 
holds the linelist table but was requested by the consortium to help 
with the iDR4 work.

Non-survey Data Release:

None.

Astrogrid deployment, VO & Data Analysis services:

Nothing.


Miscellaneous:

Nothing.
-- 
Scanned by iCritical.